OpenStudy (anonymous):

4 friends have different numbers of cows on their farms. Jorge has 3 x as many cows as Carla. Marcus has half as many cows as Jorge. Dominique has 2 less than 4 x as many cows as Carla. How many cows does each person have if they have a total of 169 cows?

OpenStudy (anonymous):

So you want to try this on your own like the last one?

OpenStudy (anonymous):

I don't know where to start

OpenStudy (anonymous):

Same like last time, lets choose letters for the names

OpenStudy (anonymous):

J=Jorge C=Carla M=Marcus D=Dominique

OpenStudy (anonymous):

Next build the relationships

OpenStudy (anonymous):

J(3)C?

OpenStudy (anonymous):

M/2=J

OpenStudy (anonymous):

Nope, M has half the number of cows as J

OpenStudy (anonymous):

So it will be M = J/2 right?

OpenStudy (anonymous):

Would it be J=C(3)?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

D-2=C(4)?

OpenStudy (anonymous):

try the last one again

OpenStudy (anonymous):

D=C(4)-2?

OpenStudy (anonymous):

yup correct!

OpenStudy (anonymous):

What's next?

OpenStudy (anonymous):

Total is 169 cows

OpenStudy (anonymous):

yup good.

OpenStudy (anonymous):

So lets write M, J and D in terms of C?

OpenStudy (anonymous):

I don't know how

OpenStudy (anonymous):

Okay M = J/2 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

But J = 3c so M = 3C/2?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

And D = 4C-2 right?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So M+J+C+D = 169 right?

OpenStudy (anonymous):

Is M=27 and C=18 and J=54 and D=70?

OpenStudy (anonymous):

Super@

OpenStudy (anonymous):

Super!

OpenStudy (anonymous):

Thanks again